OpenStudy (anonymous):

f(x)=x(8-x^2)^1/2 find where the first derivative is zero or undefined

OpenStudy (anonymous):

you got the derivative?

OpenStudy (anonymous):

no I'm stuck on it

OpenStudy (anonymous):

i take it it is \[f(x)=x\sqrt{8-x^2}\] use the product rule and chain rule to get \[f'(x)=\sqrt{8-x^2}-\frac{x^2}{\sqrt{8-x^2}}\]

OpenStudy (anonymous):

you need it step by step?

OpenStudy (anonymous):

yes, please

OpenStudy (anonymous):

k it is a product right \(x\times \sqrt{8-x^2}\) therefore you need the product rule \[(fg)'=f'g+g'f\]

OpenStudy (anonymous):

here we could say \(f(x)=x, f'(x)=1\) and \(g(x)=\sqrt{8-x^2}\)

OpenStudy (anonymous):

now we need \(g'(x)\) which requires the chain rule, or maybe just knowing that the derivative of the square root of something is the derivative of something over 2 root of something in this case you get \[g'(x)=\frac{-2x}{2\sqrt{8-x^2}}\]

OpenStudy (anonymous):

Im confused on the last message

OpenStudy (anonymous):

ok lets do a quick example what is the derivative of \(\sqrt{\sin(x)}\)? it is \[\frac{\cos(x)}{2\sqrt{\sin(x)}}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what is the derivative of \(\sqrt{x^2+3x}\) ? it is \[\frac{2x+3}{2\sqrt{x^2+3x}}\]

OpenStudy (anonymous):

and similarly the derivative of \(\sqrt{8-x^2}\) is \(\frac{-2x}{2\sqrt{8-x^2}}\) although you probably want to cancel the twos

OpenStudy (anonymous):

plug all that stuff in to here \((fg)'=f'g+g'f\) and you get \[\sqrt{8-x^2}-\frac{x^2}{\sqrt{8-x^2}}\]

OpenStudy (anonymous):

in order to find the zeros you have to actually do the subtraction to find where it is undefined is just a matter of setting \(8-x^2=0\) and solving for \(x\) pretty much in your head

OpenStudy (anonymous):

don't we add in the product rule

OpenStudy (anonymous):

that was the product rule

OpenStudy (anonymous):

\[(fg)'=f'g+g'f\] with \[f(x)=x,f'(x)=1,g(x)=\sqrt{8-x^2}, g'(x)=\frac{-x}{\sqrt{8-x^2}}\]

OpenStudy (anonymous):

plug them directly in to the formula, and see that you get (i hope) the answer i wrote above

OpenStudy (anonymous):

\[1 * \sqrt{8-x ^{2}} + 2x/2\sqrt{2-x^2} * x\] This?

OpenStudy (anonymous):

yes except you missed the minus sign

OpenStudy (anonymous):

the derivative of \(8-x^2\) is \(-2x\)

OpenStudy (anonymous):

oh ok thats where that (-) came from

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

now you have to actually do the subtraction to find the zeros

OpenStudy (anonymous):

\[\sqrt{8-x ^{2}} - x^2 / \sqrt{8-x^2}\] This is what I have now

OpenStudy (anonymous):

should I find the common denominator?

OpenStudy (anonymous):

? you only have one denominator

OpenStudy (anonymous):

duh nevermind

OpenStudy (anonymous):

what if I do then I can simplify and get rid of the first \[\sqrt{8-x^2}\]

OpenStudy (anonymous):

it will go when you subtract

OpenStudy (anonymous):

is the top where it is equal to zero and undefined on the bottom

OpenStudy (anonymous):

\[\sqrt{8-x^2}-\frac{x^2}{\sqrt{8-x^2}}\] \[=\frac{8-x^2-x^2}{\sqrt{8-x^2}}\]

OpenStudy (anonymous):

better known as \[\frac{8-2x^2}{\sqrt{8-x^2}}\]

OpenStudy (anonymous):

clear how i got that?

OpenStudy (anonymous):

yes okay the top says that the first derivative is equal to 0 when x = + or - 2 but how do I mess with the radical

OpenStudy (anonymous):

in the denominator? you don't the denominator is zero when \(8-x^2=0\)

OpenStudy (anonymous):

is that where it is undefined?

OpenStudy (anonymous):

yes, when the denominator is zero the derivative is undefined you get \(x=\pm\sqrt8\) pretty much in your head

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

not really that surprising, since those numbers are the endpoints of the domain of \(\sqrt{8-x^2}\)

OpenStudy (anonymous):

yw lotta algebra on this one huh?

OpenStudy (anonymous):

yes and I stink at algebra lol and I always thought my weakness was geometry and trig i have one more if you can help. I think its alot easier :)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

okay thanks I appreciate it quiz tomorrow on this stuff I just have to find the first derivative of \[y=3^{2x} + 3e ^{2x}-\ln(3x^2)\]

OpenStudy (anonymous):

ok lets go step by step, starting with the easy part

OpenStudy (anonymous):

sounds good to me

OpenStudy (anonymous):

the derivative of \(e^x\) is \(e^x\) therefore by the chain rule the derivative of \(e^{2x}\) is \(2e^{2x}\) making the derivative of \(e^{2x}\) as \(6e^{2x}\) so far so good?

OpenStudy (anonymous):

oops typo last line should be "making the derivative of \(3e^{2x}\) as \(6e^{2x}\)"

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

ok I got 6e^2x

OpenStudy (anonymous):

ok now how about \(3^{2x}\) or more generally \(b^x\) do you know the derivative of that?

OpenStudy (anonymous):

just b

OpenStudy (anonymous):

ooooh no!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

the derivative of \(b^x\) is \(b^x\ln(b)\)

OpenStudy (anonymous):

for example the derivative of \(10^x\) is \(10^x\ln(10)\)

OpenStudy (anonymous):

and so by the chain rule the derivative of \(3^{2x}\) is \(2\times \ln(3)\times 3^{2x}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and finally we get to the easiest one \(-\ln(3x^2)\) do you know the derivative of \(\ln(x)\) ?

OpenStudy (anonymous):

1/x

OpenStudy (anonymous):

ok so now you have a choice, you can either use the chain rule or rewrite the log, same answer in either case

OpenStudy (anonymous):

\[\ln(3x^2)=\ln(3)+2\ln(x)\] and since \(\ln(3)\) is just some number, the derivative of \(-\ln(3x^2)\) is \(-\frac{2}{x}\) that is one way

OpenStudy (anonymous):

or by the chain rule you get \[-\frac{6x}{3x^2}\] which is also \(-\frac{2}{x}\)

OpenStudy (anonymous):

ok so to clean it up a bit the first part 2 * ln3 * 3^2x how can I simplify that

OpenStudy (anonymous):

you can't really 2 and ln(3) are just numbers you can write it without the times signs, but that is all

OpenStudy (anonymous):

ok thanks you have been super helpful I really appreciate it

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

have a great night

OpenStudy (anonymous):

good luck on your quiz!